Operations Report – February 16th

Crew 275 Operations Report

16th Feb 2023
SOL: 4

Name of person filing report: Quentin ROYER
Non-nominal systems:
• Toilet
• RAM: 1 leak in the roof
• Science Dome: 3 leaks in the walls – same as yesterday
• Lower Deck carbon monoxide detector
• Propane Generator
• Oven: light of the oven non-working – heating does not seem to work either
• Spacesuit helmet
• Robotic observatory (potentially): it was opened for the whole day

Notes on non-nominal systems:
• Propane generator: power outage from 6:45 to 6:52
• Spacesuit Helmet: a zip tie is missing at the rear of one helmet, I can replace it with your permission

Spirit rover used: Yes
• Hours: 209.0
• Beginning charge: 100%
• Ending charge: 84%
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.0
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 220.9
• Beginning charge: 100%
• Ending charge: 88% – Curiosity remained ON due to a crewmember’s mistake
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.4
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Snow was removed from the rovers during the EVA, prior to departure. Curiosity remained ON after the EVA due to a crewmember’s mistake

Summary of Hab operations: Power outage from 6:45 am to 6:52 am, and the Crew reduced their energy consumption in the evening per the request of Mission Support. Routine operations were conducted at the RAM, and the Crew Engineer built the weather station which will be deployed during the next EVA.
• WATER USE: 13.39 gallons
• Water: 410.96 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Hughes Net was used during the day before the opening of COMMS for email purposes, and Starlink (MDRSCampus-Guest) is used during COMMS.

Summary of suits and radios: The radio slipped from a crewmember’s head at the very beginning of the EVA. The 3 crewmembers went back to the Hab after going through the repressurization protocol. The crewmember then wore a headband to keep the radio in place. The EVA could resume after the depressurization protocol, and Spacesuits and radios were nominal.
One helmet has a missing zip tie at its rear. I can replace it with your approval.

Summary of GreenHab operations: The Crew botanist conducted routine operations and relocated tomato plants.
• WATER USE: 12.16 gallons
• Heater: On
• Supplemental light: From 10:00 pm to 2:00 am
• Harvest: 57 g of beans

Summary of ScienceDome operations: The Commander and Crew Scientist tested some experimental protocols.
• Dual split: Heat / From 10 pm to 7 am – We suspect that the dual split did not work last night

Summary of RAM operations: The RAM was not used because of the freezing temperatures.

Summary of any observatory issues: The robotic observatory was open for the whole day.

Summary of health and safety issues: The lower deck carbon monoxide detector beeped 3 times, 3 times in a row at 21:50, it then stopped. There was a Fire Alarm in the Hab at 11:00, during the EVA. The alarm rang because the kitchen hood was not on MAX mode during cooking. The HabCom went down to the lower deck, and the EVA was not disrupted. The remaining crewmembers in the Hab managed the alarm.

Questions, concerns, and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site